# skipabove=0.5skipbelow=0.5align=center

### v1

@@ VI-E Ethical and Privacy Considerations
We propose that (i) contributors opt in to sharing reputational data, (ii) only aggregated or pseudonymized metrics (e.g., differential-privacy noise or percentile rankings) are revealed to maintainers, and (iii) contributors receive dashboards that show exactly what information will be shared. Compliance with regulations such as GDPR [83] should guide data-retention periods, user-deletion requests, and audit logging. Future prototypes should incorporate cryptographic approaches—such as zero-knowledge proofs or secure multiparty computation—to prove adherence to security metrics without exposing raw activity logs. Key directions include: Designing and evaluating privacy-preserving reputation protocols, conducting user studies to gauge contributor consent thresholds and maintainer information needs, and integrating regulatory compliance checks and automated audit trails into the prototype.
VI-F Why Focus Reputation on Cybersecurity?
@@
We have situated the ARMS approach within the context of cybersecurity. The ARMS metrics can, of course, be extended in order to infer properties of engineers other than their cybersecurity expertise. We suggest that doing so for functional properties (i.e., input-output behaviors that can be validated through testing) may be unnecessary — standard engineering practices call for code contributions to be accompanied by adequate tests, such that functional properties can be assured through reference to the test results. Not so for non-functional properties, cybersecurity as one of many. It is for such properties that reputational measures may be more useful. For example, Cramer et al. recently reported that trust and safety defect repairs in social media platforms are rarely accompanied by automated tests, in part because the validation is apparently performed through use case analysis rather than through software behavioral analysis [71]. Similarly, behaviors for regulatory compliance such as GDPR are challenging to validate [11]. Since the current state-of-the-art software validation techniques cannot automatically conclude whether a system provides non-functional properties such as security, trust-and-safety, or regulatory compliance — at least not in a cost-effective manner — we think that ARMS approaches may be a suitable complementary measure of correctness.
VII Conclusion
@@
In this paper, we explored the characteristics of the open-source software supply chain that necessitate actor-based security techniques. We advocate for the implementation of an actor reputation system to operationalize a framework for trust within the open-source ecosystem. We introduced seven security metrics for measuring reputation, alongside two evaluation methods: one for assessing actor performance and another for determining the effectiveness of these metrics in predicting reputation. Lastly, we identified future research directions, including empirical studies, system prototypes, and initiatives to maintain actor identities across ecosystems.
Appendix A Signal Metrics Definition Table
Table II: Proposed security signals, weighting factors, and evaluation metrics. Bolded security and weighting signals are derived from established frameworks (e.g., NIST, SLSA, CNCF) and prior work, respectively. The proposed metrics describe how to measure each highlighted security or weighting signal. Relevant case studies and related research are cited to justify inclusion of select metrics.
S/N. Proposed Signals & Analysis Metrics Description
SECURITY SIGNALS
1. Security vulnerability in artifact – Pull Requests/Commits Measures vulnerabilities introduced through pull requests or commits. Introduced either by the user or to a project owned by the user.
a. Time to fix/close security vulnerabilities of various severity levels [51] Time taken to address vulnerabilities detected in pull requests or commits.
b. Severity levels of security vulnerability reported [52] Evaluates the criticality of reported vulnerabilities (low, medium, high), for user’s projects.
c. Presence of vulnerability in PR/commits (not linked to external dependencies) [53] Measures severity of vulnerabilities that stem from user’s direct contributions.
2. Security vulnerability in artifact – use of vulnerable dependencies Assesses the use and introduction of vulnerable dependencies into the artifact.
a. Length of time before fix [51] Time taken to resolve vulnerable dependencies, after they have been detected (or publicized).
b. Severity levels of package’s reported vulnerability [52] Tracks severity of vulnerabilities in dependencies.
c. Number of Repos/Projects that have a vulnerable dependency Total number of projects affected by vulnerable dependencies as a percentage of users total projects.
d. Number of vulnerable dependencies/project Assesses how widespread vulnerable dependencies are within each project.
e. Number of projects with reported vulnerable dependencies Measures the overall exposure of projects to vulnerable dependencies.
3. Use of ecosystem code scanning and security analysis features Evaluates usage of ecosystem’s security tools for code scanning.
a. Status of dependabot alerts (for dependencies)[54] Monitors if security alerts for dependencies are being addressed.
b. Status of Secret sharing/Validity Checks Evaluates the management of secret-sharing mechanisms and validity checks.
c. Code scanning and Vulnerability alerts[54] Assesses if security scanning tools are actively used and vulnerability alerts managed.
d. Push Protection Status Checks if protection features are used to block pushes with exposed secrets (e.g., unscanned code).
a. Number of Projects with an integrity guarantee [27] Measures how many projects are secured with integrity verification features such as code signing.
5. Use of branch protection Checks if branch protection is enforced to prevent unauthorized changes.
a. Number of protected branches per project Evaluates how many branches within each project are safeguarded against direct commits.
b. Number of Projects with Protected branches [55, 64] Tracks how many projects enforce branch protection.
6. Use of security policies and vulnerability reporting Determines if the project has security policies and reporting mechanisms.
a. Status of private vulnerability reporting, or security policy [55] Checks if private vulnerability reporting channels and security policies are established and functional.
7. Use of automated workflows Assesses use of [55] automation in workflows to enforce security.
a. Number of Projects with Automated workflows [55] Measures how many user’s projects use automated workflows for security enforcement, such as auto-deployments or auto-tests.
REPUTATION WEIGHTAGE SIGNALS
W1. Package Usage[56]
a. Number of downloads/stars/forks Weighs the potential contributor’s ownership of useful projects to the ecosystem.
W2. Community Tenure [51]
a. Length of time contributing to project Weighs how much time user has been a contributor to other projects.
b. Length of time owning account Weighs how long user has actually owned an account
c. Contributory Strength Weighs the strength of contributions to other projects by lines of code, issues, pull requests
W3. Centrality Score[18]
a. Connections to other Actors[18] Weighs a potential contributor’s network of interactions/contributory activities.
b. Time to form connections Length of time taken to form the connections in ’a’ above.

### v2

@@ 6.6. Ethical and Privacy Considerations
We propose that (i) contributors opt in to sharing reputational data, (ii) only aggregated or pseudonymized metrics (e.g., differential-privacy noise or percentile rankings) are revealed to maintainers, and (iii) contributors receive dashboards that show exactly what information will be shared. Compliance with regulations such as GDPR (European Union, 2016) should guide data-retention periods, user-deletion requests, and audit logging. Future prototypes should incorporate cryptographic approaches—such as zero-knowledge proofs or secure multiparty computation—to prove adherence to security metrics without exposing raw activity logs. Key directions include: Designing and evaluating privacy-preserving reputation protocols, conducting user studies to gauge contributor consent thresholds and maintainer information needs, and integrating regulatory compliance checks and automated audit trails into the prototype.
6.7. Why Focus Reputation on Cybersecurity?
@@
We have situated the ARMS approach within the context of cybersecurity. The ARMS metrics can, of course, be extended in order to infer properties of engineers other than their cybersecurity expertise. We suggest that doing so for functional properties (i.e., input-output behaviors that can be validated through testing) may be unnecessary — standard engineering practices call for code contributions to be accompanied by adequate tests, such that functional properties can be assured through reference to the test results. Not so for non-functional properties, cybersecurity as one of many. It is for such properties that reputational measures may be more useful. For example, Cramer et al. recently reported that trust and safety defect repairs in social media platforms are rarely accompanied by automated tests, partly since validation is apparently performed through use case analysis rather than through software behavioral analysis (Cramer et al., 2025). Similarly, behaviors for regulatory compliance, such as GDPR are challenging to validate (Franke et al., 2024). Since the current state-of-the-art software validation techniques cannot automatically conclude whether a system provides non-functional properties such as security, trust-and-safety, or regulatory compliance, at least not in a cost-effective manner, we think that ARMS approaches may be a suitable complementary measure of correctness.
7. Conclusion
@@
In this paper, we explored the characteristics of the open-source software supply chain that necessitate actor-based security techniques. We advocate for the implementation of an actor reputation system to operationalize a framework for trust within the open-source ecosystem. We introduced seven security signals (each operationalized by concrete metrics) for estimating actor reputation, and outlined quantitative study designs to assess both signal utility and end-to-end system feasibility. Lastly, we identified future research directions, including a system prototype, quantitative validation at scale, and follow-up user studies on decision-making and deployment impacts.